# Chow and augmented Chow polynomials as evaluations of Poincaré-extended $\mathbf{a}\mathbf{b}$-indices

### v2

@@ Abstract.
We show that Chow polynomials and augmented Chow polynomials of matroids, and more generally of finite graded posets admitting $R$-labellings, are obtained as evaluations of their Poincaré-extended $\mathbf{a}\mathbf{b}$-indices. This implies in particular explicit combinatorial $\gamma$-positive expansions for both. We then evaluate this expansion to obtain a compact expansion for the braid arrangement.
@@ 1. Main results
The Chow ring $\underline{\textsf{CH}}(M)$ and the augmented Chow ring $\textsf{CH}(M)$ are important geometric-algebraic object attached to a loopless matroid $M$ that have been introduced in [FY04] and in [BHM+22], respectively. They have attracted a lot of attention in recent years, maybe most notable in their celebrated use in the proof of the log-concavity of the Whitney numbers of the second kind in [AHK18]. We follow [FMSV24] in our notations and call their Hilbert-Poincaré series $\underline{\textsf{H}}_{M}(x)$ and $\textsf{H}_{M}(x)$ the Chow polynomial and the augmented Chow polynomial.
To state the two main results, recall that the lattice of flats $\mathcal{L}(M)$ of the matroid $M$ is geometric and thus admits an $R$-labeling $\lambda$, i.e., an edge-labeling $\lambda$ such that each interval in $\mathcal{L}(M)$ contains a unique maximal chain along which $\lambda$ is increasing [Sta74]. To a maximal chain =$\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in $\mathcal{L}(M)$ associate the sequence $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ of edge labels $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$ for $n={\sf rank}(M)$ being the rank of $M$.
@@
$\underline{\textsf{H}}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,,$
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{1}<\lambda_{2}$ and also that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$.
@@
$\textsf{H}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,,$
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$.
@@
The Chow polynomial $\underline{\textsf{H}}_{M}(x)$ and the augmented Chow polynomial $\textsf{H}_{M}(x)$ are $\gamma$-positive. They are in particular nonnegative, palindromic and unimodal.
To the best of our knowledge, the only known proof of the $\gamma$-positivity appears in [FMSV24] and relies on semi-small decompositions for the Chow and the augmented Chow rings [BHM+22] or, equivalently, on the Kähler package for these rings [AHK18]. Hence, Theorems 1.1, LABEL: and 1.2 may be viewed as a Hodge-theory free proof to the $\gamma$-positivity and also of unimodality of Chow and augmented Chow polynomials of geometric lattices. After proving the main theorems in Section 2, we evaluate the expansions in the case of the braid arrangement in Section 3.
@@
In forthcoming work by Ferroni–Matherne–Vecchi, the definitions of the Chow and augmented Chow polynomials are generalized to finite graded posets with $\hat{0}$ and $\hat{1}$ [FMV24], see Definition 2.1. Using this definition for general finite graded posets, our proofs hold for all such posets admitting $R$-labelings. We in particular obtain non-negativity, palindromicity, unimodality, and $\gamma$-positivity in this generality.
To prove the main results, we show for finite graded posets admitting $R$-labelings that the Chow and the augmented Chow polynomials are obtained as evaluations of the Poincaré-extended $\mathbf{a}\mathbf{b}$-index defined and studied by the author and Dorpalen-Barry–Maglione in [DMS23]. It turns out that the Chow polynomial, but not the augmented Chow polynomial, is an evaluation of the coarse flag Hilbert-Poincaré series as defined and studied by Maglione–Voll in [MV23], which itself is also an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. As the Poincaré-extended $\mathbf{a}\mathbf{b}$-indices of finite graded poset admitting $R$-labelings have explicit descriptions using the labeling along maximal chains, we then deduce Theorems 1.1, LABEL: and 1.2.
@@
We show in Theorem 2.6 that the Chow polynomial is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. This evaluation is conjectured to be real-rooted [FS23, Conjecture 8.18]. It is shown in [DMS23, Corollary 2.23] that the numerator polynomial of the coarse flag Hilbert-Poincaré series is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index and asked in [KM23, Question 1.6] for which matroids a further evaluation is real-rooted. Also the chain polynomial of the lattice of flats is yet another evaluation of the extended $\mathbf{a}\mathbf{b}$-index which is conjectured in [AKE23, Conjecture 1.2] to be real-rooted. This suggests the following question.
@@ Acknowledgment
This work was conducted at the Mathematisches Forschungsinstitut Oberwolfach in June 2024 during the workshop “Arrangements, Matroids and Logarithmic Vector Fields”. We very much thank the institute for their hospitality and the organizers Takuro Abe, Graham Denham, Eva-Maria Feichtner, and Gerhard Röhrle for setting up this wonderful and inspiring meeting. The results in this paper were obtained after attending the talk by Luis Ferroni entitled “Chow polynomials of posets”. The author would very much like to thank Luis and his collaborators Chris Eur, Jacob Matherne, and Lorenzo Vecchi, and as well Elena Hoster and Lukas Kühne for inspiring discussions concerning Chow polynomials and their combinatorial interpretations, and for comments on a preliminary version of this paper.
@@ 2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index and its evaluations
Let $P$ be a finite graded poset of rank $n$. That is, $P$ is a finite poset with unique minimum element $\hat{0}$ of rank 00 and unique maximum element $\hat{1}$ of rank $n$ such that ${\sf rank}(F)$ is equal to the length of any maximal chain from $\hat{0}$ to $F$. Its Möbius function $\mu$ is given by $\mu(F,F)=1$ for all $F\in P$ and $\mu(F,G)=-\sum_{F\leq H<G}\mu(F,H)$ for all $F<G$. Its Poincaré polynomial is
@@
A (not necessary maximal) chain $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ in $P$ is an ordered set of pairwise comparable elements. We assume throughout this paper that all chains end in the maximum element $\mathcal{C}_{k+1}=\hat{1}$. A chain is maximal if $k=n$ and we usually write $\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in this case in order to clearly distinguish between general chains ending in $\hat{1}$ and maximal chains. The chain Poincaré polynomial and the reduced chain characteristic polynomial of such a chain $\mathcal{C}$ are
@@
where $[\mathcal{C}_{i},\mathcal{C}_{i+1}]\subseteq P$ denotes the interval between two consecutive elements in the chain. In particular,
@@ 2.1. The Chow and the augmented Chow polynomials
The following description of the Chow and augmented Chow polynomials of a matroid is taken from [FMSV24, Theorem 1.3 & 1.4]. We present it in the generality of general finite graded posets as introduced by [FMV24].
Definition 2.1 (Ferroni–Matherne–Vecchi).
@@ 2.2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index
The following description and properties of the extended $\mathbf{a}\mathbf{b}$-index are taken from [DMS23]. Throughout this section, let $P$ be a finite graded poset admitting an $R$-labeling, i.e., an edge-labeling $\lambda:\mathcal{E}(P)\rightarrow\mathbb{N}_{+}$ such that each interval in $P$ contains a unique maximal chain along which $\lambda$ is (weakly) increasing.
@@
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{Z}[y]\langle\mathbf{a},\mathbf{b}\rangle\,,$
@@
(2) ${\text{\scriptsize\sf ex}}\Psi_{P}(0,\mathbf{a},\mathbf{b})=\Psi_{P}(\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$
@@
$\mathbf{b}\cdot\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{N}[y]\langle\mathbf{a},\mathbf{b}\rangle\,,$
@@
Observe that the factor $\mathbf{b}$ on the left-hand side comes from the assumption that $\mathcal{C}_{1}=\hat{0}$ and thus, every $\mathbf{a},\mathbf{b}$-polynomial ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ starts with the variable $\mathbf{b}$ as $0={\sf rank}(\mathcal{C}_{1})$. For later reference, we set
(3) ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}\,.$
We next present two of the main result of [DMS23], from which we then deduce the proposed evaluation to obtain the Chow and the augmented Chow polynomials. As in Section 1, we write $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ for the labeling along a maximal chain =$\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ with $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$. For a subset $E\subseteq\{1,\dots,n\}$, we moreover set $\lambda_{\mathcal{F},E}=(\lambda^{\prime}_{0},\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{n})$ with $\lambda^{\prime}_{0}=0$ and $\lambda^{\prime}_{i}=-\lambda_{i}$ if $i\in E$ and $\lambda^{\prime}_{i}=\lambda_{i}$ if $i\notin E$. We in addition define ${\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})=u_{0}\cdots u_{n-1}$ to be the monomial in $\mathbf{a},\mathbf{b}$ with
$u_{i}=\begin{cases}\mathbf{a}&\text{if }\lambda^{\prime}_{i}\leq\lambda^{\prime}_{i+1}\,,\\ \mathbf{b}&\text{if }\lambda^{\prime}_{i}>\lambda^{\prime}_{i+1}\,,\\ \end{cases}$
@@
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{F},E}y^{\#E}\cdot{\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})$
@@
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\omega\big{(}\Psi_{P}(\mathbf{a},\mathbf{b})\big{)}$

### v3

@@ Abstract.
We show that Chow polynomials and augmented Chow polynomials of matroids, and more generally of finite graded posets admitting $R$-labelings, are obtained as evaluations of their Poincaré-extended $\mathbf{a}\mathbf{b}$-indices. This implies in particular explicit combinatorial $\gamma$-positive expansions for both, providing the first proof of the $\gamma$-positivity not relying on the Kähler package for the Chow ring. We then evaluate this expansion to obtain an explicit closed formula for the braid arrangement.
@@ 1. Main results
The Chow ring and the augmented Chow ring are important algebro-geometric objects attached to a loopless matroid that have been introduced in [FY04] and in [BHM+22], respectively. They have attracted a lot of attention in recent years, maybe most notable in their celebrated use in the proof of the log-concavity of the Whitney numbers of the first kind in [AHK18]. We follow [FMSV24] in our notations and call their Hilbert-Poincaré series $\underline{\textsf{H}}_{M}(x)$ and $\textsf{H}_{M}(x)$ the Chow polynomial and the augmented Chow polynomial.
Recall that the lattice of flats $\mathcal{L}(M)$ of a matroid $M$ is geometric and thus admits an $R$-labeling $\lambda$, i.e., an edge-labeling $\lambda$ such that each interval in $\mathcal{L}(M)$ contains a unique maximal chain along which $\lambda$ is (weakly) increasing [Sta74]. For our purposes, we assume $\lambda$ to have nonnegative integer values. To a maximal chain $\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in $\mathcal{L}(M)$ associate the sequence $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ of edge labels $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$ for $n={\sf rank}(M)$. We moreover set ${\sf Des}(\mathcal{F})=\{i\mid\lambda_{i}>\lambda_{i+1}\}$ and $\mathsf{des}(\mathcal{F})=|{\sf Des}(\mathcal{F})|$, and we call ${\sf Des}(\mathcal{F})$ isolated [^1] This property has been called good in [FMV24] and stable in [BV25]. if $i\in{\sf Des}(\mathcal{F})$ implies $i+1\notin{\sf Des}(\mathcal{F})$.
The following two theorems are the main results of this paper.
@@
$\underline{\textsf{H}}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\mathcal{F})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\mathcal{F})}\,$,
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that ${\sf Des}(\mathcal{F})$ is isolated and $1\notin{\sf Des}(\mathcal{F})$.
@@
$\textsf{H}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\mathcal{F})}(x+1)^{n-2\ \!\!\mathsf{des}(\mathcal{F})}\,$,
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that ${\sf Des}(\mathcal{F})$ is isolated.
@@
The Chow polynomial $\underline{\textsf{H}}_{M}(x)$ and the augmented Chow polynomial $\textsf{H}_{M}(x)$ are $\gamma$-positive. They are, in particular, nonnegative, palindromic and unimodal.
To the best of our knowledge, the only known proof of the $\gamma$-positivity appears in [FMSV24] and relies on semi-small decompositions for the Chow and the augmented Chow rings [BHM+22] or, equivalently, on the Kähler package for these rings [AHK18]. Hence, Theorems 1.1, LABEL: and 1.2 may be viewed as a Hodge-theory free proof to the $\gamma$-positivity and also of unimodality of Chow and augmented Chow polynomials of geometric lattices. We prove Theorems 1.1, LABEL: and 1.2 in Section 2. In Section 3, we then use those results to provide new explicit closed formulas for the (augmented) Chow polynomial in the case of the braid arrangement.
@@
In work by Ferroni–Matherne–Vecchi, the definitions of the Chow and augmented Chow polynomials are generalized to finite graded posets with $\hat{0}$ and $\hat{1}$ [FMV24], see Definition 2.1. Using this definition for general finite graded posets, our proofs hold for all such posets admitting $R$-labelings. We, in particular, obtain non-negativity, palindromicity, unimodality, and $\gamma$-positivity in this level of generality.
To prove the main results, we show for finite graded posets admitting $R$-labelings that the Chow and the augmented Chow polynomials are obtained as evaluations of the Poincaré-extended $\mathbf{a}\mathbf{b}$-index defined and studied by the author and Dorpalen-Barry–Maglione in [DBMS25]. It turns out that the Chow polynomial, but not the augmented Chow polynomial, is an evaluation of the coarse flag Hilbert-Poincaré series as defined and studied by Maglione–Voll in [MV24], which itself is also an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. As the Poincaré-extended $\mathbf{a}\mathbf{b}$-indices of finite graded poset admitting $R$-labelings have explicit descriptions using the labeling along maximal chains, we then deduce Theorems 1.1, LABEL: and 1.2.
@@
In Theorem 2.6, we show that the Chow polynomial is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. This evaluation is conjectured to be real-rooted [FS23, Conjecture 8.18]. It is shown in [DBMS25, Corollary 2.23] that the numerator polynomial of the coarse flag Hilbert-Poincaré series is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index and asked in [KM23, Question 1.6] for which matroids a further evaluation is real-rooted. Also the chain polynomial of the lattice of flats is yet another evaluation of the extended $\mathbf{a}\mathbf{b}$-index which is conjectured in [AKE23, Conjecture 1.2] to be real-rooted. This suggests the following question.
@@ Acknowledgment
This work was conducted at the Mathematisches Forschungsinstitut Oberwolfach in June 2024 during the workshop “Arrangements, Matroids and Logarithmic Vector Fields”. We very much thank the institute for their hospitality and the organizers Takuro Abe, Graham Denham, Eva-Maria Feichtner, and Gerhard Röhrle for setting up this wonderful and inspiring meeting. The results in this paper were obtained after attending the talk by Luis Ferroni entitled “Chow polynomials of posets”. The author would very much like to thank Luis and his collaborators Chris Eur, Jacob Matherne, and Lorenzo Vecchi, and as well Elena Hoster and Lukas Kühne for inspiring discussions concerning Chow polynomials and their combinatorial interpretations, and for comments on a preliminary version of this paper. He also thanks three anonymous referees for their thoughtful comments that helped to improve the presentation of this paper.
@@ 2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index and its evaluations
Let $P$ be a finite graded poset of rank $n$. That is, $P$ is a finite poset with unique minimum element $\hat{0}$ and unique maximum element $\hat{1}$ of rank $n$ such that ${\sf rank}(F)$ is equal to the length of any maximal chain from $\hat{0}$ to $F$. Its Möbius function $\mu$ is given by $\mu(F,F)=1$ for all $F\in P$ and $\mu(F,G)=-\sum_{F\leq H<G}\mu(F,H)$ for all $F<G$. Its Poincaré polynomial is
@@
A chain $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ in $P$ is an ordered set of pairwise comparable elements. We assume throughout this paper that all chains end in the maximum element $\mathcal{C}_{k+1}=\hat{1}$. A chain is maximal if $k=n$ and we usually write $\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in this case in order to clearly distinguish between general chains ending in $\hat{1}$ and maximal chains. The chain Poincaré polynomial and the reduced chain characteristic polynomial of such a chain $\mathcal{C}$ are
@@
where $\mathcal{C}_{i},\mathcal{C}_{i+1}]\subseteq P$ denotes the interval between two consecutive elements in the chain. In particular,
@@ 2.1. The Chow and the augmented Chow polynomials
The following description of the Chow and augmented Chow polynomials of a matroid is taken from [FMSV24, Theorem 1.3 & 1.4]. We present it for general finite graded posets as introduced by [FMV24].
Definition 2.1 ([FMV24, Definition 3.3]).
@@ 2.2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index
The following description and properties of the extended $\mathbf{a}\mathbf{b}$-index are taken from [DBMS25]. Throughout this section, let $P$ be a finite graded poset admitting an $R$-labeling, i.e., an edge-labeling $\lambda:\mathcal{E}(P)\rightarrow\mathbb{N}_{+}$ such that each interval in $P$ contains a unique maximal chain along which $\lambda$ is (weakly) increasing.
@@
${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{Z}[y]\langle\mathbf{a},\mathbf{b}\rangle\,$,
@@
(2) ${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(0,\mathbf{a},\mathbf{b})=\Psi_{P}(\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$
@@
$\mathbf{b}\cdot\iota\big({\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big)=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{N}[y]\langle\mathbf{a},\mathbf{b}\rangle\,$,
@@
Observe that the factor $\mathbf{b}$ on the left-hand side comes from the assumption that $\mathcal{C}_{1}=\hat{0}$ and thus every $\mathbf{a},\mathbf{b}$-polynomial ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ starts with the variable $\mathbf{b}$ as $0={\sf rank}(\mathcal{C}_{1})$. For later reference, we set
(3) ${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big({\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big)\,.$
We next present two of the main result of [DBMS25], from which we then deduce the proposed evaluation to obtain the Chow and the augmented Chow polynomials. As in Section 1, we write $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ for the labeling along a maximal chain $\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ with $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$. For a subset $E\subseteq\{1,\dots,n\}$, we moreover set $\lambda_{\mathcal{F},E}=(\lambda^{\prime}_{0},\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{n})$ with $\lambda^{\prime}_{0}=0$ and $\lambda^{\prime}_{i}=-\lambda_{i}$ if $i\in E$ and $\lambda^{\prime}_{i}=\lambda_{i}$ if $i\notin E$. We in addition define ${\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})=u_{0}\cdots u_{n-1}$ to be the monomial in $\mathbf{a},\mathbf{b}$ with
[…]
@@
${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{F},E}y^{\#E}\cdot{\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})$
@@
${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\omega\big(\Psi_{P}(\mathbf{a},\mathbf{b})\big)$